Logiweb codex of append

Up Help append, append(∗, ∗),

append

[append →^{pyk} “append”]

append(∗, ∗)

[append(x, y) →^{val} y : : x] [append(x, y) →^{tex} “append(#1., #2.)”] The pyk compiler, version 0.grue.20050502+ by Klaus Grue GRD-2005-05-26.UTC:20:37:04.090427 = MJD-53516.TAI:20:37:36.090427 = LGT-4623856656090427e-6